OpenStudy (anonymous):

Help please word problem what sign i should use how much hotter is the melting point of zinc than that of sodium sodium is 98 and zinc is 420

OpenStudy (anonymous):

what sign should i use to do this problem

OpenStudy (anonymous):

divided multi add or substract ??

OpenStudy (anonymous):

?? dont know

OpenStudy (anonymous):

Well, if Zinc is 420 and sodium 98, then the difference between the two is 420-98= 322.

OpenStudy (anonymous):

ok so u subtract can u help with a few more ??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how much warmer is venus than pluto... pluto is 480 and venus is -230 we are doing intergers

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

yeah, subtract the two getting 710

OpenStudy (anonymous):

u sure its not -710

OpenStudy (anonymous):

Sorry, yes it is. Are you sure those values are correct though? Venus should be hotter than pluto...

OpenStudy (anonymous):

i got 250

OpenStudy (anonymous):

when i subtracted them both

OpenStudy (anonymous):

oh integers

OpenStudy (anonymous):

Venus' temperature is plus minus 460 and about -300 or so... BTW, -230-480 = -710. Not 480-230... that is actually adding them as 480+(-230) is what you did

OpenStudy (anonymous):

They are all integers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok next one what is the difference btw the melting points of aluminum and copper ... alumin is 659 and copper is 1083

OpenStudy (anonymous):

rspencer u still with me ?

OpenStudy (anonymous):

yes. Subtract 659 from 1083 to get 424

OpenStudy (anonymous):

I have more 1. how much hotter is the boiling pont of water than propane ... propane is boils at -44 and water is 212 2. what is the difference btw the two freezing points ... gas -150 and water 32

OpenStudy (anonymous):

I hope u having fun with this lol

OpenStudy (anonymous):

This is just simple subtraction... do you need me for this?

OpenStudy (anonymous):

no but for both of them

OpenStudy (anonymous):

and what about the neg

OpenStudy (anonymous):

so 1. is 68 - or no

OpenStudy (anonymous):

Take one value and subtract the other. If you want to check, see which is bigger and whether or not it should be negative. Remember, a bigger number minus a smaller number is positive but the other way round for a smaller number minus a bigger number

OpenStudy (anonymous):

and 2. is 118

OpenStudy (anonymous):

so 1. is not a neg but num 2. is a neg

OpenStudy (anonymous):

Number 2: yes. Number 1: no. 212-(-44) = 212+44 = 256

OpenStudy (anonymous):

so 2. is -118

OpenStudy (anonymous):

No, you had number 2 right

OpenStudy (anonymous):

ok two more if u dont mind I have more but i dont think u want them.

OpenStudy (anonymous):

Last 2

OpenStudy (anonymous):

ok 118

OpenStudy (anonymous):

What is the difference in temperature ... Mercury 400 c day side and dark side is -180 c .

OpenStudy (anonymous):

400-(-180) = 400+180=580

OpenStudy (anonymous):

I kinda gotta go. Check you later.

OpenStudy (anonymous):

and lastly which freezes first ? how much colder does it have to be for both to freeze? ... oxygen is -218.79 c and hydrogen is -252.8 c

OpenStudy (anonymous):

ok bi thanks a lot